# High Performance Distributed Control Law for Large-Scale Linear Systems: A Partitioned Distributed Observer Approach

### v3

@@ IV-B Performance of partitioned distributed observer
$\dot{V}_{i}\leq$ $-\gamma\theta^{n}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}+\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\bar{\sigma}(A_{ii}+A_{ii}^{T})\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\theta^{n-1}\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\left\|\eta_{l\star}\right\|\left\|\eta_{\star i}^{(p)}\right\|$
$+2\theta^{n-1}\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(\|A_{i}\|W_{T_{1}}+\|B_{i}\|\kappa_{i}\right)|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|$
$\leq$ $-\gamma\theta^{n}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}+\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\bar{\sigma}(A_{ii}+A_{ii}^{T})\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\theta^{n-1}\bar{\sigma}(P_{i})\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{l\star}\right\|^{2}$
$+\theta^{n-1}\bar{\sigma}(P_{i})\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\theta^{n-1}\bar{\sigma}(P_{i})\left(\|A_{i}\|W_{T_{1}}+\|B_{i}\|\kappa_{i}\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|.$
@@
$\dot{V}\leq$ $-\gamma\theta^{n}\underline{\lambda}_{\mathcal{O}_{p}}\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}+\lambda_{A}\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\lambda_{P}\theta^{n-1}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{l\star}\right\|^{2}$
$+\lambda_{P}\theta^{n-1}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\lambda_{P}\theta^{n-1}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|$
$\leq$ $-\gamma\theta^{n}\underline{\lambda}_{\mathcal{O}_{p}}\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\theta^{n-1}\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\|\eta_{i\star}\|^{2}$
$+\theta^{n-1}\left(\lambda_{A}+\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\theta^{n-1}\lambda_{P}\bar{\lambda}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|,$
where $\kappa=\max_{i=1,\ldots,N}\{\kappa_{i}\}$. It is noticed that Lemma 4 leads to
$\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\|\eta_{\star i}^{(p)}\|^{2}=\sum_{i=1}^{N}\|\eta_{i\star}\|^{2},$
and
$\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\|\eta_{\star i}^{(p)}\|\leq\sqrt{2}\|\eta\|,$
where $\eta=col\{\eta_{i\star},~{}i=1,\dots,N\}$. Then, by denoting $c(\theta)=\gamma\theta\underline{\lambda}_{\mathcal{O}_{p}}-\lambda_{A}/\theta^{n-1}-2\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)$, and $c_{K}=2\lambda_{P}\bar{\lambda}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)$, we can further obtain
$\dot{V}\leq$ $-c(\theta)\theta^{n-1}\|\eta\|^{2}+c_{K}\theta^{n-1}\|\eta\|.$ (26)
Condition 3) indicates that $c(\theta)>0$, and $c(\theta)$ is obviously monotonically increasing and radially unbounded with respect to $\theta$.
@@
This section has proved that error dynamics of the partitioned distributed observer can converge to an any small invariant set before the given time $T_{1}$. Due to the inability to analyze the performance of distributed observers and closed-loop systems separately, Theorem 1 relies on assumptions about $T_{1}$. In Theorem 3 of the next subsection, we will provide a joint analysis of the observer and controller, where the assumption about $T_{1}$ will be removed. Next, we will further elaborate on some details and supplementary explanations of Theorem 1 in several remarks.
@@
The design methods for the parameters of the partitioned distributed observer are detailed in Theorem 1. According to Theorem 1, all parameters except for $\gamma$ only rely on the information of each individual agent and do not require any other information. Designing $\gamma$ involves the information of system matrix, input matrix, and network topology. However, $\gamma$ is actually just a sufficiently large constant. Therefore, in practical usage, it is often avoided to use global information by designing $\gamma$ as an adaptive parameter.
@@
Assumption 1 requires the observability of $(C_{i},A_{ii})$. Note that $\bar{C}_{i}$ and $\bar{A}_{ii}$ are obtained by a similarity transformation from $C_{i}$ and $A_{i}$. Thus [9, Lemma 3.1] shows that $(\bar{C}_{i},\bar{A}_{ii})$ is also observable. It means $\bar{H}_{i}$ given in condition 1) of Theorem 1 can be obtained through pole configuration.
Remark 7
Fusion estimation and model mismatch in partitioned distributed observer design make the stability analysis of its error dynamics (10)–(11) a real challenge because they not only bring more complex error dynamics but also derive two distinct compact error forms ($e_{\star i}^{(p)}$ and $e_{i\star}$). In this subsection, Lemma 4 demonstrates the strict mathematical relationship between these two compact error forms. Then, Theorem 1 develops the two-layer Lyapunov analysis method, which ingeniously transforms the above mathematical relationship into the results that can be used in the traditional Lyapunov stability analysis. The proof of Lemma 4 and the design of the two-layer Lyapunov function ($V=\sum_{i=1}^{N}V_{i}$) are the keys to successfully proving the error dynamic stability of the partitioned distributed observer.
@@ IV-C Closed-loop performance under the distributed control law
At the beginning, one obtains by centralized control law (12) that the closed-loop system with (12) is
@@
Consider the closed-loop system (31) as well as its communication network and physical network subject to Assumption 1 and 2. If the error $\|e\|$ of the partitioned distributed observer (8)–(9) stays in $\Omega_{e}$ with $t\geq 0$, then $\|x\|$ can converge to an invariant
@@
This is the actual control law adopted by the closed-loop system in this paper, and also the source of output information $y_{i}$ in the partitioned distributed observer (8)–(9). Let $\bar{U}=col\{\bar{u}_{i}-u_{i},~{}i=1,\ldots,N\}$ and obtain the compact form
@@
Theorem 3 is of primary importance for this paper. It combines the conclusions of Theorem 1 and Theorem 2, and obtains the final result: the error system and the closed-loop system of the partitioned distributed observer can converge to any small invariant sets. Furthermore, the finite time conditions involved in Theorem 1 are eliminated by the analysis in Theorem 3. In fact, the final conclusion of the theorem only depends on the controllability and observability of the system as well as the connectivity of the communication network and does not depend on any other assumptions. These conclusions also indicate that, compared with the traditional distributed observer, the performance of the partitioned distributed observer in this paper has almost no loss, but the dimension of observers on each agent is greatly reduced. These results make the distributed observer more practical in large-scale system problems.
@@ IV-D The ability of performance recovery
Since $K$ is chosen so that $A+BK$ is a Hurwitz matrix, the system (29) is stable. Hence, there exists $T_{2}>0$ such that $\|x_{c}(t)\|\leq\varepsilon/2$ for all $t>T_{2}$. Besides, according to Theorem 3, there is $M_{1}>0$ so that $24c_{K}\|QB\|\|K\|/c(\theta)c_{2}<\varepsilon/$ 2 for all $\theta>M_{1}$. Then, there is a constant $T_{3}>0$ such that $\|x_{r}(t)\|<\varepsilon/2$ for all $\theta>M_{1}$ and $t>T_{3}$. Subsequently, we have
@@
Finally, we consider the scenario where $t\in(T(\theta),T]$. We define $F_{c}(t,x)=(A+BK)x$ and $F_{r}(t,x,e)=(A+BK)x+B\bar{U}$, where $\bar{U}$ is an implicit function of $e$ in light of Theorem 2 and 3. Since $(T(\theta),T]$ is a finite time interval, the conditions of the famous theorem (See in any monograph of “Ordinary Differential Equation”) named “The Continuous Dependence Theorem of Solutions of Differential Equations on Initial Values and Parameters” are fulfilled. It indicates that there is a constant $\delta>0$ such that $\|x_{c}(t)-x_{r}(t)\|<\varepsilon$ if the parameter $e$ in $F_{r}(t,x,e)$ satisfying $\|e\|<\delta$. Bearing in mind the conclusion of Theorem 1 and 3, there exists $M_{3}>0$ such that $\|e\|<\delta^{\prime}(\theta)<\delta$ for all $\theta>M_{3}$.
@@
Up to now, the goals of this paper have been fully realized. After partition, the dimension of each local observer is far less than that of the interconnected system, but the distributed control law designed based on the partition distributed observer can still ensure that the controlled system can approximate the centralized control performance arbitrarily.
Remark 9
In the process of implementing the partitioned distributed observer and the distributed-observer-based distributed control law, the high-gain parameter $\theta$ plays an important role. However, it should be noted that the existence of high gain makes this method unable to deal directly with measurement noise. When measurement noise exists, there is a trade-off between the steady-state error (caused by measurement noise), the steady-state error (caused by model mismatch), and the difference between the distributed control performance and the centralized control performance. The smaller the latter two, the larger the former. Therefore, when both measurement noise and model mismatch exist, finding an optimal $\theta$ to consider these three indicators simultaneously is a topic that is worth further research in the future.
@@ V Simulation
To illustrate the effectiveness of the developed methods, this section will be divided into three parts. Firstly, we describe the simulation system. Then, the effectiveness of the network partitioning method will be demonstrated. Finally, we show the validity of the partitioned distributed observer and the distributed-observer-based distributed control law.
@@ V-A System formulation
In this paper, we randomly select $\tau_{P_{i}}$ within $[0.012,0.018]$, and randomly select $k_{P_{i}}$ within $[1\times 10^{-15},10\times 10^{-15}]$. Besides, we choose $V_{i}=110{\rm V}$ for all $i=1,\ldots,N$.
V-B Effectiveness of partition method
@@
Each agent needs to estimate the whole system’s states if the classical distributed observer method [12, 13, 15, 16] is used. Therefore, the dimension of observers on each agent is $94$. In contrast, it can be greatly reduced if the partitioned distributed observer developed in this paper is employed. See in Figure 8; the purple bar shows that, after partitioning, the maximum observer dimension of all agents is only $20$ (reduced by $78.72\%$), which is much lower than that of the traditional method (yellow bar). In addition, after partitioning, the minimum observer dimension on each agent is only $4$ (reduced by $95.7\%$), and the average value is only $10.8083$ (reduced by $88.51\%$). Both of them are much lower than the value of the yellow bar. By the way, the merging partition process in Algorithm 1 in this paper can also effectively reduce the dimension of observers on each agent. By comparing the blue with purple bars, we see that the maximum and average values of the local observer dimensions after the partition merging (purple bar) are smaller than those before the partition merging (blue bar). Besides, through MATLAB simulation, the partition process of this example takes $0.017$ seconds.
@@
Figure 9: Physical and communication network with 100 nodes.
Figure 10: Variation of dynamic process performance with $\theta$.
To further illustrate the effectiveness of the partition algorithm, this paper also selects a system with $100$ nodes. Its physical network and communication network are shown in Figure 9. In this example, the similarity between the two networks is $78.49\%$. As seen in Figure 10, compared to the traditional distributed observer, the maximum dimension of observers on each agent in the proposed partitioned distributed observer is reduced by $84.00\%$, and the average dimension is reduced by $92.46\%$. In this example, the time consumption of the partition algorithm by MATLAB is $0.07$ seconds.
In conclusion, the partition algorithm proposed in this paper can reduce the average local observer dimension by about $90\%$ when the similarity between the physical network and the communication network is about $80\%$. Furthermore, the time consumption of the developed algorithm is very small.

### v4

@@ IV-B Performance of cover-based distributed observer
$\dot{V}_{i}\leq$ $-2\gamma\theta^{n-1}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}+\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{A_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\left\|\eta_{l\star}\right\|\left\|\eta_{\star i}^{(p)}\right\|$
$+2\bar{\sigma}(P_{i})\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left(\|A_{i}\|W_{T_{1}}+\|B_{i}\|\kappa_{i}\right)|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|$
$\leq$ $-2\gamma\theta^{n-1}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}+\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\lambda_{A_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\bar{\sigma}(P_{i})\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{l\star}\right\|^{2}$
$+\bar{\sigma}(P_{i})\left(\|A_{i}\|+\rho_{i}\|B_{i}\|\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\bar{\sigma}(P_{i})\left(\|A_{i}\|W_{T_{1}}+\|B_{i}\|\kappa_{i}\right)\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|.$
@@
$\dot{V}\leq$ $\left(-2\gamma\theta^{n-1}\underline{\lambda}_{\mathcal{O}_{p}}+\lambda_{A_{i}}\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\lambda_{P}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{l\star}\right\|^{2}$
$+\lambda_{P}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\sum_{l\in\mathcal{O}_{p}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\lambda_{P}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}|\mathcal{O}_{p}|\left\|\eta_{\star i}^{(p)}\right\|$
$\leq$ $\left(-2\gamma\theta^{n-1}\underline{\lambda}_{\mathcal{O}_{p}}+\lambda_{A_{i}}\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\|\eta_{i\star}\|^{2}$
$+\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|^{2}$
$+2\lambda_{P}\bar{\lambda}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)\sum_{i=1}^{N}\sum_{\mathcal{O}_{p}\in\mathcal{P}_{i}}\left\|\eta_{\star i}^{(p)}\right\|,$
where $\kappa=\max_{i=1,\ldots,N}\{\kappa_{i}\}$. Note that Lemma 4 indicates ∑_i=1^N∑_O_p∈P_i∥η_⋆i^(p)∥^2=∑_i=1^N∥η_i⋆∥^2, and ∑_i=1^N∑_O_p∈P_i∥η_⋆i^(p)∥≤2∥η∥, where $\eta=col\{\eta_{i\star},~i=1,\dots,N\}$. Then, by denoting $c(\theta)=2\gamma\theta^{n-1}\underline{\lambda}_{\mathcal{O}_{p}}-\lambda_{A_{i}}-2\lambda_{P}\bar{\lambda}\left(\|A\|+\rho\|B\|\right)$, and $c_{K}=2\lambda_{P}\bar{\lambda}\left(\|A\|W_{T_{1}}+\|B\|\kappa\right)$, we can further obtain
$\dot{V}\leq$ $-c(\theta)\|\eta\|^{2}+c_{K}\|\eta\|.$ (26)
Condition 3) indicates that $c(\theta)>0$, and $c(\theta)$ is obviously monotonically increasing and radially unbounded with respect to $\theta^{n-1}$.
@@
This section has proved that error dynamics of the cover-based distributed observer can converge to an any small invariant set before the given time $T_{1}$. Due to the inability to analyze the performance of distributed observers and closed-loop systems separately, Theorem 1 relies on assumptions about $T_{1}$. In Theorem 3 of the next subsection, we will provide a joint analysis of the observer and controller, where the assumption about $T_{1}$ will be removed. Next, we will further elaborate on some details and supplementary explanations of Theorem 1 in several remarks.
@@
The design methods for the parameters of the cover-based distributed observer are detailed in Theorem 1. According to Theorem 1, all parameters except for $\gamma$ only rely on the information of each individual agent. Though designing $\gamma$ involves the information of system matrix, input matrix, and network topology, $\gamma$ is actually just a sufficiently large constant. Therefore, in practical usage, it is often avoided to use global information by designing $\gamma$ as an adaptive parameter [7, 38].
@@
Fusion estimation and model mismatch in cover-based distributed observer design make the stability analysis of its error dynamics (10)–(11) a real challenge because they not only bring more complex error dynamics but also derive two distinct compact error forms ($e_{\star i}^{(p)}$ and $e_{i\star}$). In this subsection, Lemma 4 demonstrates the strict mathematical relationship between these two compact error forms. Then, Theorem 1 develops the two-layer Lyapunov analysis method, which ingeniously transforms the above mathematical relationship into the results that can be used in the traditional Lyapunov stability analysis. The proof of Lemma 4 and the design of the two-layer Lyapunov function ($V=\sum_{i=1}^{N}V_{i}$) are the keys to successfully proving the error dynamic stability of the cover-based distributed observer.
@@ IV-C Closed-loop performance under the distributed control law
At the beginning, one obtains that the closed-loop system with centralized control law (12) is
@@
Consider the closed-loop system (31) as well as its communication network and physical network subject to Assumption 1 and 2. If the error $\|e\|$ of the cover-based distributed observer (8)–(9) stays in $\Omega_{e}$ with $t\geq 0$, then $\|x\|$ can converge to an invariant
@@
This is the actual control law adopted by the closed-loop system in this paper, and also the source of output information $y_{i}$ in the cover-based distributed observer (8)–(9). Let $\bar{U}=col\{\bar{u}_{i}-u_{i},~i=1,\ldots,N\}$ and obtain the compact form
@@
Theorem 3 is of primary importance for this paper. It combines the conclusions of Theorem 1 and Theorem 2, and obtains the final result: the error system and the closed-loop system of the cover-based distributed observer can converge to any small invariant sets. Furthermore, the finite time conditions involved in Theorem 1 are eliminated by the analysis in Theorem 3. In fact, the final conclusion of the theorem only depends on the controllability and observability of the system as well as the connectivity of the communication network and does not depend on any other assumptions. These conclusions also indicate that, compared with the traditional distributed observer, the performance of the cover-based distributed observer in this paper has almost no loss, but the dimension of observers on each agent is greatly reduced. These results make the distributed observer more practical in large-scale system problems.
@@ IV-D The ability of performance recovery
Since $K$ is chosen so that $A+BK$ is a Hurwitz matrix, the system (29) is stable. Hence, there exists $T_{2}>0$ such that $\|x_{c}(t)\|\leq\varepsilon/2$ for all $t>T_{2}$. Besides, according to Theorem 3, there is $M_{1}>0$ so that $4c_{K}\|QB\|\|K\|/c(\theta)c_{2}<\varepsilon/2$ for all $\theta>M_{1}$. Then, there is a constant $T_{3}>0$ such that $\|x_{r}(t)\|<\varepsilon/2$ for all $\theta>M_{1}$ and $t>T_{3}$. Subsequently, we have
@@
Finally, we consider the scenario where $t\in(T(\theta),T]$. We define $F_{c}(t,x)=(A+BK)x$ and $F_{r}(t,x,e)=(A+BK)x+B\bar{U}$, where $\bar{U}$ is an implicit function of $e$ in light of Theorem 2 and 3. Since $(T(\theta),T]$ is a finite time interval, the conditions of the famous theorem named “The Continuous Dependence Theorem of Solutions of Differential Equations on Initial Values and Parameters” [39] are fulfilled. It indicates that there is a constant $\delta>0$ such that $\|x_{c}(t)-x_{r}(t)\|<\varepsilon$ if the parameter $e$ in $F_{r}(t,x,e)$ satisfying $\|e\|<\delta$. Bearing in mind the conclusion of Theorem 1 and 3, there exists $M_{3}>0$ such that $\|e\|<\delta^{\prime}(\theta)<\delta$ for all $\theta>M_{3}$.
@@
Up to now, the goals of this paper have been fully realized. After find coverage, the dimension of each local observer is far less than that of the interconnected system, but the distributed control law designed based on the cover-based distributed observer can still ensure that the controlled system can approximate the centralized control performance arbitrarily.
@@ V Simulation
To illustrate the effectiveness of the developed methods, this section will be divided into three parts. Firstly, we describe the simulation system. Then, the effectiveness of the coverage solving method will be demonstrated. Finally, we show the validity of the cover-based distributed observer and the distributed-observer-based distributed control law.
@@ V-A System formulation
In this paper, we randomly select $\tau_{P_{i}}$ within $[0.012,0.018]$, and randomly select $k_{P_{i}}$ within $1\times 10^{-15},10\times 10^{-15}]$. Besides, we choose $V_{i}=110{\rm V}$ for all $i=1,\ldots,N$.
V-B Effectiveness of coverage solving method
@@
Each agent needs to estimate the whole system’s states if the classical distributed observer method [6, 7, 9, 10] is used. Therefore, the dimension of observers on each agent is $94$. In contrast, it can be greatly reduced if the cover-based distributed observer developed in this paper is employed. See in Figure 8; the purple bar shows that, after constructing coverage, the maximum observer dimension of all agents is only $20$ (reduced by $78.72\%$), which is much lower than that of the traditional method (yellow bar). In addition, after constructing coverage, the minimum observer dimension on each agent is only $4$ (reduced by $95.7\%$), and the average value is only $10.8083$ (reduced by $88.51\%$). Both of them are much lower than the value of the yellow bar. By the way, the merging process in Algorithm 1 in this paper can also effectively reduce the dimension of observers on each agent. By comparing the blue with purple bars, we see that the maximum and average values of the local observer dimensions after merging cover sets (purple bar) are smaller than those before merging sets (blue bar).